OpenStudy (anonymous):

Evaluate m – n when m = –1.5 and n = –7.7. A. –9.2 B. –6.2 C. 6.2 D. 9.2

OpenStudy (anonymous):

D mabe

OpenStudy (anonymous):

What is -1.5-(-7.7) = ?

TheSmartOne (thesmartone):

Hi

OpenStudy (anonymous):

See - times - is +

OpenStudy (rippeh):

C

OpenStudy (anonymous):

hi smarty

TheSmartOne (thesmartone):

Just substitute the values given to you into the equation

OpenStudy (anonymous):

C

OpenStudy (anonymous):

So u ll get -1.5 + 7.7=

OpenStudy (anonymous):

Which equals 6.2

TheSmartOne (thesmartone):

Note: -1.5 + 7.7 = 7.7 - 1.5